# The Complexity of Proper Homotopy Equivalence of Graphs

### v1

@@ 1. Introduction
1.1. Mapping class groups
The mapping class group of a surface is the group of homeomorphisms of that surface considered up to homotopy. The group of outer automorphisms of a free group can be realized as the group of homotopy equivalences of a graph whose fundamental group is that free group. These two groups of symmetries are closely related and classical objects of study in geometric group theory. Traditionally, the study of these objects focused on finite-type surfaces, those with finitely generated fundamental groups, and finite graphs. The corresponding groups of symmetries are then finitely generated groups. Recently, there has been a considerable rise in interest in infinite-type surfaces, those with infinitely-generated fundamental group, and their mapping class groups. In parallel, Algom-Kfir and Bestvina introduced the mapping class group of an infinite graph, a geometric infinite-type analog of $\operatorname{Out}(F_{n})$, in [AB21].
The mapping class group of a graph is defined to be the group of proper homotopy equivalences, up to proper homotopy. Thus, graphs which are themselves proper homotopy equivalent have the same (ie isomorphic) mapping class groups, just as homeomorphic surfaces have the same mapping class groups. So, it is natural to ask the parallel questions:
@@ 1.1. Mapping class groups
1.2. Classification of definable equivalence relations
To make the above questions precise, we use the terminology of complexity of equivalence relations from descriptive set theory. In general, given an equivalence relation $E$ on a Polish space $X$, we say that $E$ is at most as complicated as the equivalence relation $F$ on Polish space $Y$, and we write $E\leq_{B}F$, if there is a Borel reduction $\phi:X\to Y$ in the sense that
@@ 1.2. Classification of definable equivalence relations
Intuitively, existence of such a reduction says that to check if $x$ and $y$ are $E$-related, we can pass the points to $Y$ via $\phi$ and check if they are $F$-related. Thus, $E$ cannot be any “harder” of a classification problem than $F$. The condition that the reduction must be Borel is to ensure that the reduction is constructive (otherwise, we could always use the axiom of choice to map each $E$-class to a unique $F$-class). If we have $E\leq_{B}F$ and $F\leq_{B}E$, we say that $E$ and $F$ are bireducible and we write $E\equiv_{B}F$. Studying the relative complexities of equivalence relations on Polish spaces, up to bireducibility, has been a major recent focus of descriptive set theory. The following are some important points in the hierarchy of analytic equivalence relations in increasing order of complexity, see also fig. 1.
  • All orbit equivalence relations of countable groups reduce to the shift action of $\mathbb{F}_{2}$, the free group on two generators, on $2^{\mathbb{F}_{2}}$. This is due to Dougherty, Jackson, and Kechris in [DJK94]. This orbit equivalence relation is sometimes called $\mathbb{E}_{\infty}$.
  • All orbit equivalence relations of non-archimedean group actions reduce to a universal $S_{\infty}$ action called $C_{\infty}$. This is due to Becker and Kechris in [BK96]. If $E\equiv_{B}C_{\infty}$, then $E$ is sometimes called Borel complete in the literature.
  • All orbit equivalence relations of Polish group actions reduce to the homeomorphism relation of compact metric spaces. This is due to Zielinski in [Zie16].
@@
Figure 1. A partial illustration of analytic equivalence relations up to bireducibility. Below the dotted line are the Borel equivalence relations, where each equivalence class is a Borel set.
1.3. Results for proper homotopy equivalence of graphs
We first realize the set of locally finite graphs as a Polish space, in Section 2.2, which makes proper homotopy equivalence an analytic equivalence relation. Our main result is the following.
@@ 1.3. Results for proper homotopy equivalence of graphs
Proper homotopy equivalence on the space of locally finite graphs is bireducible with $C_{\infty}$. That is, the equivalence relation $\operatorname{\thicksim_{\text{PHE}}}$ is Borel complete.
We get the same complexity result when considering the subspaces of graphs of uniform or bounded degree. It should be noted that while being bireducible with $C_{\infty}$ is often reffered to as being “Borel complete,” the equivalence relations we are considering are not Borel equivalence relations. Our proof relies heavily on the result of Camerlo and Gao in [CG01] that the equivalence relation of homeomorphism on closed subsets of the Cantor set is bireducible with $C_{\infty}$.
Figure 2. The Loch Ness Monster graph (top) and surface (bottom).
The graph formed by taking a ray and attaching infinitely many loops is often called the Loch Ness Monster graph (see fig. 2). This was after the corresponding surface, constructed from the plane by attaching handles along the positive real axis was named the Loch Ness Monster surface. Our second result shows that the equivalence classes of this graph is generic among infinite graphs.
@@
The proper homotopy equivalence class of the Loch Ness Monster graph is a dense $G_{\delta}$ set in the space of locally finite, infinite graphs.
1.4. Results for homeomorphism of surfaces
In order to realize orientable surfaces as a Polish space, we endow them with a pants decomposition. Here a pants decomposition can mean a collection of simple closed curves on the surface up to isotopy so that the complementary components are $d$-holed spheres, for fixed, bounded, or finite $d$. See Section 2.3 for precise definitions of the spaces under consideration with examples.
@@ 1.4. Results for homeomorphism of surfaces
A version of the above corollary (where the result is the same, but the Polish space of surfaces is different) was proven by Bergfalk and Smythe in their upcoming preprint.
The analogous corollary to 1.2 for surfaces says that the homeomorphism class of the Loch Ness Monster surface is a dense $G_{\delta}$ set in the space of noncompact surfaces with a pants decomposition. We interpret this result as saying the following: If you randomly glue pairs of pants together to build a noncompact surface, that surface will almost surely be homeomorphic to a Loch Ness Monster surface. This holds true if you allow your building blocks to include annuli or pants with arbitrarily many legs.
1.5. Outline
In Section 2, we give some background information about the spaces we are working in and the tools we use. In Section 3 we prove 1.1 and 1.3 in the space of $3$-regular graphs. We show how to extend the proofs given in Section 3 to other spaces of graphs and surfaces in Section 4.
@@ Acknowledgements
The authors would like to thank Denis Osin for suggesting this project along with a starting plan of attack. Many thanks to Christian Rosendal for helpful conversations. Much appreciation to Lucy Kristoffersen for providing a new perspective. The first author was gratefully supported by NSF DMS–2303365 (MSPRF).
2. Background
@@ 2.1. Graphs
Figure 3. An example of a graph where $v_{1}$ has two self loops, and there are multiple edges between $v_{2}$ and $v_{3}$. The degree of $v_{1}$ is $5$, the degrees of $v_{2}$ and $v_{3}$ are $4$, and the degree of $v_{4}$ is $1$.
We realize a graph as a topological space by building the CW-complex with 00-cells corresponding to the vertex set of $\Gamma$ and $1$-cells corresponding to the edge set of $\Gamma$. The $1$-cells are then glued to the 00-cells by the incidence relation. The graph has a natural metric, called the path metric, which assigns each edge length one, so that adjacent vertices are distance one from each other. A graph is connected if it is connected as a CW complex; for the rest of the paper we will assume all graphs are connected. Note that connected, locally finite, infinite graphs necessarily have infinite but countable vertex and edge sets.
The fundamental group of a (connected) graph is always a free group, so we define the rank of a graph to be $\operatorname{rk}(\Gamma):=\operatorname{rk}(\pi_{1}(\Gamma))$. It is an exercise in algebraic topology to show that two graphs are homotopy equivalent exactly when they have the same rank. Because we will focus on infinite graphs, we will consider the equivalence relation of proper homotopy equivalence. Recall that a function between topological spaces is proper if the preimage of compact sets is compact. A proper homotopy equivalence is a continuous map $f:X\rightarrow Y$ which is proper with proper homotopy inverse $g:Y\rightarrow X$, so that both $f\circ g$ and $g\circ f$ are properly homotopic to identity maps. We study the equivalence relation of proper homotopy equivalence (PHE), and we will write $\Gamma\operatorname{\thicksim_{\text{PHE}}}\Delta$ if $\Gamma$ and $\Delta$ are PHE. The classification of locally finite, infinite graphs up to proper homotopy equivalence was done by Ayala, Dominguez, Márquez, and Quintero in [ADMQ90]. In order to understand the classification, we need to understand the end space of a graph.
@@
where $K$ are compact subsets of ${\Gamma}$.
The end space will always be Hausdorff, totally disconnected, and compact, so a Stone space. Because graphs are Hausdorff, locally compact, and $\sigma$-compact, we can find a countable sequence of compact sets $K_{i}$ that obtain $E(\Gamma)$. Let $e\in E({\Gamma})$, and define the $i$th neighborhood of $e$, denoted $N_{i}(e)$, to be the component of $\Gamma\setminus K_{i}$ which contains $e$. We say that the end $e$ is accumulated by loops if $\operatorname{rk}(N_{i}(e))=\infty$ for all $i$. The set of ends of ${\Gamma}$ accumulated by loops is a closed subspace of $E(\Gamma)$, denoted by $E_{\ell}({\Gamma})$. Together $(E({\Gamma}),E_{\ell}({\Gamma}))$ are called the endspace pair of ${\Gamma}$. Endspace pairs $(E(\Gamma),E_{\ell}({\Gamma}))$ and $(E(\Delta),E_{\ell}(\Delta))$ are isomorphic if there is a homeomorphism $\phi:E(\Gamma)\rightarrow E(\Delta)$ with $\phi(E_{\ell}(\Gamma))=E_{\ell}(\Delta)$. We are now ready to state the proper homotopy equivalence classification theorem for graphs.
Theorem 2.2 ([ADMQ90, Theorem 2.7]).
@@
[…]
Figure 4. Examples of graphs with different end spaces. Each of these graphs have different PHE types.
@@ 2.2. Spaces of Graphs
Two graphs will be distance zero exactly when there is an isomorphism between them fixing the basepoint, so the metric topology is Haurdorff. The $\epsilon$-balls of the basepoints of any two graphs in $\mathcal{G}_{k}$ will be isomorphic for all ${\epsilon}<\frac{1}{2}$. Thus, the distance function is always finite valued and $\mathcal{G}_{k}$ is finite diameter. The sets
@@
form a clopen basis for the topology. The basis is clopen because the sets $\mathcal{U}_{({\Gamma},r)}$ and $\mathcal{U}_{({\Gamma},r+{\epsilon})}$ will agree when $[r,r+{\epsilon}]$ does not contain an integer or half-integer. Note that it is again very important that our graphs be connected, otherwise this topology will not be Hausdorff.
@@
The space $\mathcal{G}_{1}$ is empty, and $\mathcal{G}_{2}$ is a singleton, corresponding to the graph which is homeomorphic to a line. The space $\mathcal{G}_{3}$ is our first infinite, non-discrete, Polish space, as are $\mathcal{G}_{k}$ for all $k\geq 4$. However, none of these spaces contain all PHE-types of infinite graphs. For example, there is only one $k$-regular tree for each $k$, and for $k\geq 3$ these all have end space homeomorphic to the Cantor set. So, we also consider the spaces $\mathcal{G}_{\leq k}$, which consists of infinite graphs where each vertex has degree less than or equal to $k$. We topologize the space in the exact same way as $\mathcal{G}_{k}$ above, allowing $r=0$, so that graphs with different degree basepoints are distance $1$ from each other. Each space $\mathcal{G}_{\leq k}$ for $k\geq 3$ contains every proper homotopy equivalence type of graph. To see this, start with the construction of standard models in Definition 2.5 of [AB21] and possibly apply proper homotopy equivalences as in fig. 5 to decrease the degree of any vertex. The final space we consider is $\mathcal{G}_{<\infty}$, which is the space of all locally finite, infinite graphs, again topologized in the same way. We will use the notation $\mathcal{G}_{\star}$ to denote any of the previously described spaces of graphs.
@@
The degree of a vertex is not invariant under proper homotopy equivalence, so while we study the equivalence relation of PHE, note that the homotopies only necessarily live in $\mathcal{G}_{<\infty}$.

### v2

@@ 1. Introduction
The mapping class group of a surface is the group of homeomorphisms of that surface considered up to homotopy. The group of outer automorphisms of a free group can be realized as the group of homotopy equivalences of a graph whose fundamental group is that free group. These two groups of symmetries are closely related and classical objects of study in geometric group theory. Traditionally, the study of these objects focused on finite-type surfaces, those with finitely generated fundamental groups, and finite graphs. The corresponding groups of symmetries are then finitely generated groups. Recently, there has been a considerable rise in interest in infinite-type surfaces, those with infinitely-generated fundamental group, and their mapping class groups. In parallel, Algom-Kfir and Bestvina introduced the mapping class group of an infinite graph, a geometric infinite-type analog of $\operatorname{Out}(F_{n})$, in [AB2021].
The mapping class group of a graph is defined to be the group of proper homotopy equivalences, up to proper homotopy. Thus, graphs which are themselves proper homotopy equivalent have isomorphic mapping class groups, just as homeomorphic surfaces have isomorphic mapping class groups. So, it is natural to ask the parallel questions:
@@
To make the above questions precise, we use the terminology of complexity of equivalence relations from descriptive set theory. In general, given an equivalence relation $E$ on a standard Borel space $X$, we say that $E$ is at most as complicated as the equivalence relation $F$ on standard Borel space $Y$, and we write $E\leq_{B}F$, if there is a Borel reduction $\phi:X\to Y$ in the sense that
@@
Intuitively, existence of such a reduction says that to check if $x$ and $y$ are $E$-related, we can pass the points to $Y$ via $\phi$ and check if they are $F$-related. Thus, $E$ cannot be any “harder” of a classification problem than $F$. The condition that the reduction must be Borel is to ensure that the reduction is constructive. Otherwise, assuming that there are at least as many $F$-classes as there are $E$-classes, we could always use the axiom of choice to map each $E$-class to a unique $F$-class. If we have $E\leq_{B}F$ and $F\leq_{B}E$, we say that $E$ and $F$ are bireducible and we write $E\equiv_{B}F$. Studying the relative complexities of equivalence relations up to bireducibility, has been a major recent focus of descriptive set theory.
Example 1.1.
The following are some important points in the hierarchy of analytic equivalence relations in increasing order of complexity.
  • All orbit equivalence relations of countable groups reduce to the shift action of $\mathbb{F}_{2}$, the free group on two generators, on $2^{\mathbb{F}_{2}}$. This is due to Dougherty, Jackson, and Kechris in [DJK94].
  • All orbit equivalence relations of non-archimedean group actions reduce to a universal $S_{\infty}$ action called $C_{\infty}$. This is due to Becker and Kechris in [BK96]. This is the complexity of isomorphism of countable connected graphs.
  • All orbit equivalence relations of Polish group actions reduce to the homeomorphism relation of compact metric spaces. This is due to Zielinski in [Z16].
@@
We first realize the set of locally finite graphs as a standard Borel space in two different ways, in Section 2.3, which makes proper homotopy equivalence an analytic equivalence relation. Our main result is the following.
@@
Proper homotopy equivalence on the space of locally finite graphs is bireducible with $C_{\infty}$.
We get the same complexity result when considering subspaces of graphs of uniform or bounded degree. Our proof relies heavily on the result of Camerlo and Gao in [CG01] that the equivalence relation of homeomorphism on closed subsets of the Cantor set is bireducible with $C_{\infty}$. The lower bound in our complexity result follows a similar line of reasoning as a response on MathOverflow by Conley, where he outlines a reduction of homeomorphism of open sets in the plane to $C_{\infty}$ using the result of Camerlo and Gao [Con].
Other equivalence relations at this level include isomorphism of countable graphs ([H93]) and isomorphism of countable linear orderings ([FS89]). These are each examples of isomorphisms of countable structures, which Borel reduce to $C_{\infty}$. It should be noted that while being bireducible with $C_{\infty}$ is often referred to as being “Borel complete,” the equivalence relations we are considering are not Borel equivalence relations.
Figure 1. Representations of the Loch Ness monster graph (top) and surface (bottom).
The graph formed by taking a ray and attaching infinitely many loops is often called the Loch Ness monster graph, see fig. 1. This was after the corresponding surface, constructed from the plane by attaching handles along the positive real axis was named the Loch Ness monster surface. Our second result shows that the equivalence classes of this graph is generic among infinite graphs.
@@
The proper homotopy equivalence class of the Loch Ness monster graph is comeager in the space of locally finite, infinite graphs.
We note that while the PHE equivalence class of the Loch Ness monster graph is comeager, this does not come from a comeager isomorphism class. In fact, in the spaces of graphs considered, every isomorphism class is countable.
In order to realize orientable surfaces as a standard Borel space, we endow them with a pants decomposition. Here a pants decomposition can mean a collection of simple closed curves on the surface up to isotopy so that the complementary components are $d$-holed spheres, for fixed, bounded, or finite $d$. See Section 4.1 for precise definitions with examples.
@@
In upcoming work, Bergfalk and Smythe establish another standard Borel space of surfaces and show that the homeomorphism relation is also bireducible with $C_{\infty}$, [BS].
The analogous corollary to 1.3 for surfaces says that the homeomorphism class of the Loch Ness monster surface is a comeager set in the space of noncompact surfaces with a pants decomposition. We interpret this result as saying the following: If you randomly glue pairs of pants together to build a noncompact surface, that surface will almost surely be homeomorphic to a Loch Ness monster surface. This holds true if you allow your building blocks to include annuli or pants with arbitrarily many legs.
Outline
In Section 2, we give some background information about the spaces we are working in and the tools we use. In Section 3 we first prove the aforementioned results (1.2 and 1.4) for 3-regular graphs. We then extend the results to other spaces of graphs, including bounded-degree graphs and locally finite graphs. Finally, we show how to extend these proofs to spaces of surfaces in Section 4.
@@ Acknowledgements
The authors would like to thank Denis Osin for suggesting this project along with a starting plan of attack. Many thanks to Christian Rosendal, Forte Shinko, and Iian Smythe for helpful conversations. Much appreciation to Lucy Kristoffersen for providing a new perspective. The first author was gratefully supported by NSF DMS–2303365 (MSPRF).
2. Preliminaries
@@ 2.1. Graphs
$v_{1}$$v_{2}$$v_{3}$$v_{4}$
Figure 2. An example of a graph where $v_{1}$ has two self loops, and there are multiple edges between $v_{2}$ and $v_{3}$. The degree of $v_{1}$ is $5$, the degrees of $v_{2}$ and $v_{3}$ are 44, and the degree of $v_{4}$ is $1$.
We realize a graph as a topological space by building the CW-complex with 0-cells corresponding to the vertex set of $\Gamma$ and $1$-cells corresponding to the edge set of $\Gamma$. The $1$-cells are then glued to the 0-cells by the incidence relation. The graph has a natural metric, called the path metric, which assigns each edge length one, so that adjacent vertices are distance one from each other. A graph is connected if it is connected as a CW complex; for the rest of the paper we will assume all graphs are connected. Note that connected, locally finite, infinite graphs necessarily have infinite but countable vertex and edge sets.
The fundamental group of a graph is always a free group, so we define the rank of a graph to be $\operatorname{rk}(\Gamma):=\operatorname{rk}(\pi_{1}(\Gamma))$. Recall that a continuous function between topological spaces, $f:X\rightarrow Y$, is a homotopy equivalence if there is a continuous $g:Y\rightarrow X$ such that $fg$ is homotopic to the identity map, $\operatorname{id}_{Y}$, and $gf\simeq\operatorname{id}_{X}$. A function between topological spaces is proper if the preimage of compact sets is compact. Thus, a proper homotopy equivalence is a continuous map $f:X\rightarrow Y$ which is proper with proper homotopy inverse $g:Y\rightarrow X$, so that both $f\circ g$ and $g\circ f$ are properly homotopic to identity maps. It is an exercise in algebraic topology to show that two graphs are homotopy equivalent exactly when they have the same rank. Because we will focus on infinite graphs, we will consider the equivalence relation of proper homotopy equivalence (PHE), and we will write $\Gamma\operatorname{\thicksim_{\text{PHE}}}\Delta$ if $\Gamma$ and $\Delta$ are PHE.
2.2. End spaces of graphs
The classification of locally finite, infinite graphs up to proper homotopy equivalence was done by Ayala, Dominguez, Márquez, and Quintero in [ayala1990proper]. In order to understand the classification, we need to define the end space of a graph.
@@ 2.2. End spaces of graphs
where $\pi_{0}$ denotes the zeroth homotopy set and $K$ ranges over compact subsets of $\Gamma$}.
Because we consider locally finite graphs, they are proper and $\sigma$-compact. That is, closed metric balls are compact, and a countable collection of compact sets cover the graphs. Specifically, for any $\Gamma\in\mathcal{G}_{\star}$ the collection of closed integer radius metric balls about the basepoint, $\{\overline{B_{n}({\Gamma})}\}_{n\in\mathbb{N}}$, covers $\Gamma$ and can be used to obtain $E(\Gamma)$. Call these closed balls $B_{n}$. The transition map $\rho_{n}:\pi_{0}(\Gamma\setminus B_{n+1})\twoheadrightarrow\pi_{0}(\Gamma\setminus B_{n})$ is induced from the inclusion map $\Gamma\setminus B_{n+1}\hookrightarrow\Gamma\setminus B_{n}$.
Recall that the inverse limit topology is the subspace topology in the product space $\prod_{\mathbb{N}}\pi_{0}(\Gamma\setminus B_{n})$. Note that there are finitely many complimentary components to each $B_{n}$, so each coordinate in the product space is discrete. Let $U$ be a complementary component of $B_{n}$, then $U$ defines a basis element for the topology on the end space by
$U^{*}=\{e\in E({\Gamma})\mid\rho_{n}(e)=U\}.$
If one visualizes an end as a “way to walk to infinity” in the graph $\Gamma$}, then the projection of $e$ to the $n$th coordinate is the complimentary component of $B_{n}$ which contains the tail of any path to $e$.
One can check that the topology described above is Hausdorff, totally disconnected, and compact by Tychonoff’s theorem. The basis given above is countable, because there are only finitely many complimentary components to each $B_{n}$, so $E(\Gamma)$ is second countable. Geometric topologists may be familiar with the refrain, “As such, the endspace is homeomorphic to a closed subset of the Cantor set.” In fact, we will later identify the endspace of a graph with a closed subset of Baire space in a Borel way, as a crucial step to proving our main theorem.
Let $e\in E({\Gamma})$, and define the $i$th neighborhood of $e$, denoted $N_{i}(e)$, to be the component of $\Gamma\setminus B_{i}$ which contains $e$. We say that the end $e$ is accumulated by loops if $\operatorname{rk}(N_{i}(e))>0$ for all $i$, or equivalently, if $\operatorname{rk}(N_{i}(e))=\infty$ for all $i$. The set of ends of $\Gamma$} accumulated by loops is a closed subspace of $E(\Gamma)$, denoted by $E_{\ell}({\Gamma})$. Together $(E({\Gamma}),E_{\ell}({\Gamma}))$ are called the endspace pair of $\Gamma$}. Endspace pairs $(E(\Gamma),E_{\ell}({\Gamma}))$ and $(E(\Delta),E_{\ell}(\Delta))$ are isomorphic if there is a homeomorphism $\phi:E(\Gamma)\rightarrow E(\Delta)$ with $\phi(E_{\ell}(\Gamma))=E_{\ell}(\Delta)$. We are now ready to state the proper homotopy equivalence classification theorem for graphs.
Theorem 2.2 ([ayala1990proper, Theorem 2.7]).
@@
Figure 3 shows three graphs which are not proper homotopy equivalent to each other. The first graph is rank $3$ and has end space homeomorphic to $(\{\ast\},\emptyset)$, a singleton and the empty set. The second has infinite rank and end space homeomorphic to $(\{\frac{1}{n}\}\cup\{0\},\{0\})$, with the subspace topology from $\mathbb{R}$. The third graph has a Cantor set of ends, with a clopen subset (so itself homeomorphic to a Cantor set) of ends accumulated by loops.
Figure 3. Examples of graphs with different PHE types.
@@ 2.3. Spaces of graphs
We realize locally finite graphs as standard Borel spaces in two different ways. Then we show that $\operatorname{\thicksim_{\text{PHE}}}$ on these spaces are Borel bireducible.
2.3.1. Based Graphs
@@
Two graphs will be distance zero exactly when there is an isomorphism between them fixing the basepoint, making the topology Hausdorff. The $\epsilon$-balls of the basepoints of any two graphs in $\mathcal{G}_{k}$ will be isomorphic for all ${\epsilon}<\frac{1}{2}$. Thus, the distance function is always finite valued and $\mathcal{G}_{k}$ is finite diameter. The sets
@@
form a clopen basis for the topology. The basis is clopen because the sets $\mathcal{U}_{({\Gamma},r)}$ and $\mathcal{U}_{({\Gamma},r+{\epsilon})}$ will agree when $r,r+{\epsilon}]$ does not contain an integer or half-integer. Note that it is again very important that our graphs be connected, otherwise this topology will not be Hausdorff.
@@
The space $\mathcal{G}_{1}$ is empty, and $\mathcal{G}_{2}$ is a singleton, corresponding to the graph which is homeomorphic to a line. The space $\mathcal{G}_{3}$ is our first infinite, non-discrete, Polish space, as are $\mathcal{G}_{k}$ for all $k\geq 4$. However, none of these spaces contain all PHE-types of infinite graphs. For example, there is only one $k$-regular tree for each $k$, and for $k\geq 3$ these all have end space homeomorphic to the Cantor set. So, we also consider the spaces $\mathcal{G}_{\leq k}$, which consists of infinite graphs where each vertex has degree less than or equal to $k$. We topologize the space in the exact same way as $\mathcal{G}_{k}$ above, allowing $r=0$, so that graphs with different degree basepoints are distance $1$ from each other. Each space $\mathcal{G}_{\leq k}$ for $k\geq 3$ contains every proper homotopy equivalence type of graph. To see this, start with the construction of standard models in Definition 2.5 of [AB2021] and possibly apply proper homotopy equivalences as in fig. 4 to decrease the degree of any vertex. The final space we consider is $\mathcal{G}_{<\infty}$, which is the space of all locally finite, infinite graphs, again topologized in the same way. We will use the notation $\mathcal{G}_{\star}$ to denote any of the previously described spaces of graphs.
@@
The degree of a vertex is not invariant under proper homotopy equivalence, so while we study the equivalence relation of PHE, note that the homotopies do not live in $\mathcal{G}_{\star}\times[0,1]$.
$\sim$